Two friends start running in different directions from the same place. One heads south, while the other heads east. After 3 hours, they are 30 miles apart. How fast is each walking on average?
Answer by ewatrrr(24785)

D = rt
Assuming the same average rate
Applying the Pythagorean Theorem
(r(3hr))^2 + (r(3hr))^2 = 30^2
18r^2 = 900
r^2 = 50
r = 5√2 mph